Tip Top Corp. produces a product that requires 6 standard hours per unit at a standard hourly rate of $14 per hour. If 3,000 units required 17,500 hours at an hourly rate of $14.56 per hour, what is the direct labor (a) rate variance, (b) time variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

a. Direct labor rate variance $
b. Direct labor time variance $
c. Direct labor cost variance $
Answer #1

Direct Labor rate variance = (Standard rate-actual rate)actual hour = (14-14.56)*17500 = 9800 Unfavorable

Direct labor time variance = (Standard hour-actual hour)Standard rate = (3000*6-17500)*14 = -7000 Favorable

Direct labor cost variance = (3000*6*14)-(17500*14.56) = 2800 Unfavorable